Kyle had 36 books in his locker. Some were library books and the rest were telephone books. the number of library books and telephone books combined is equal to twice the number of textbooks. the number of textbooks and telephone books combined is equal to 3 times the number of library books. how many of each type of book were in Kyle's locker?
Answer by stanbon(75887)

I'll set up the equations and let you do the solving.
Let number of library be "a"
Let number of telephone be "b"
Let number of textbook be "c"
Then:
a + b + c = 36
c + b = 3a
a + b = 2c
Rearrange to get :
a + b + c = 36
3a -b -c = 0
a + b -2c = 0
Answer is a=9, b=15, 12
Cheers,
Stan H.